OpenStudy (anonymous):

What is the X-value of the solution to this system of equations? 3x+2y=6 2x+y=2

OpenStudy (paki):

@KenDallBenDall please follow my steps...

OpenStudy (anonymous):

Thank you :-)

OpenStudy (paki):

or... by simply multiply the first equation (3x+2y=6 ) by 2... and the second equation (2x+y=2) by 3... we will get by doing so... 6x+4y=12... equation A... 6x+3y=6....equation B... now just subtract equation B from A... you will get y=6...

OpenStudy (paki):

@KenDallBenDall pleasure dear...